# Real-world models for multiple term structures: a unifying HJM semimartingale framework

### v2

@@ 3.5. An example: the minimal market model
We present a simple example of a model which falls into the real-world HJM setup developed above, while not admitting a risk-neutral probability. The example corresponds to an extension of the minimal market model (MMM), see [46, Chapter 13] and [13, Section 3.2]. Denoting by $X^{0}=\exp(\int_{0}^{\cdot}r_{s}{\rm d}s)$ the numéraire, we recall that in the stylized MMM the $X^{0}$-discounted growth-optimal portfolio (GOP) process $\bar{X}^{*}=(\bar{X}^{*}_{t})_{t\geq 0}$ satisfies the SDE
@@
with a one-dimensional Brownian process $W$ and where the drift $\alpha^{*}$ is modeled as a function of the form $\alpha^{*}(t)=\alpha_{0}\exp(\eta t)$, for all $t\geq 0$. The model is parameterized by a positive initial value $\alpha_{0}>0$ and a constant net growth rate $\eta>0$. We let $\{S^{i};i\in I\}$ be a family of strictly positive processes representing the spot processes. Similarly to [44, Section 3], we impose the following assumption. [^4] We recall that two processes $X=(X_{t})_{t\geq 0}$ and $Y=(Y_{t})_{t\geq 0}$ on $(\Omega,\mathbb{F},\mathbb{P})$ are said to be $\mathbb{F}$-conditionally independent if $\mathbb{E}[f(X_{T})g(Y_{T})|\mathcal{F}_{t}]=\mathbb{E}[f(X_{T})|\mathcal{F}_{t}]\mathbb{E}[g(Y_{T})|\mathcal{F}_{t}],\qquad\text{ for all }0\leq t\leq T<+\infty,$ for all measurable functions $f,g:\mathbb{R}\to\mathbb{R}_{+}$. This property is obviously satisfied if one of the two processes is deterministic.
@@
It is well-known that the MMM does not admit a risk-neutral probability (see [46, Section 13.3]). Therefore, prices are computed by relying on the real-world pricing formula (see [46, Section 10.4]). In particular, the price $B^{0}(t,T)$ of a riskless bond is given by
@@
The example described in this section illustrates a generic way of constructing real-world models satisfying market viability, without requiring risk-neutral probabilities. One can start from a strictly positive local martingale $Z$, a numéraire $X^{0}$ and a family $\{S^{i}:i\in I\}$ of strictly positive spot processes. Generalizing (3.23) and (3.25), one can then specify riskless and risky bond prices as
(3.26) $B^{i}(t,T):=\mathbb{E}\left[\frac{Z_{T}}{Z_{t}}\frac{X^{0}_{t}}{X^{0}_{T}}\frac{S^{i}_{T}}{S^{i}_{t}}\bigg{|}\mathcal{F}_{t}\right],\qquad\text{ for all}0\leq t\leq T<+\infty\text{ and }i\in I_{0},$
where $S^{0}\equiv 1$. By construction, $Z$ is a LMD and, therefore, the conditions of Theorem 3.4 are satisfied. Moreover, all term structures are fairly priced by $Z$, in the sense of Definition 3.8.
@@ 4.1. A general existence and uniqueness result for SPDEs
Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered probability space as introduced in Section 3.1 and supporting a $d$-dimensional Brownian motion $W$ and a homogeneous Poisson random measure $\mu({\rm d}t,{\rm d}x)$ on +×$\mathbb{R}_{+}\times E$ with compensator $\nu({\rm d}t,{\rm d}x)=F({\rm d}x){\rm d}t$, where $E$ is a locally compact space equipped with its Borel sigma-field $\mathcal{B}(E)$ and $F$ is a sigma-finite measure on $(E,\mathcal{B}(E))$. The compensated random measure is denoted by $\tilde{\mu}({\rm d}t,{\rm d}x):=\mu({\rm d}t,{\rm d}x)-F({\rm d}x){\rm d}t$.
@@
Let $\overline{W}$ be the $\mathbb{R}^{d+1}$-valued semimartingale given by $\overline{W}_{t}:=(t,W_{t})$, for all $t\in\mathbb{R}_{+}$. Let $r\in\mathbb{R}_{+}$ be arbitrary and let $A$ be a control process for $\overline{W}$ in the sense of [43, Definition 23.13], which exists by [43, Theorem 23.11]. Since $L^{r}$ is locally bounded, we can define the process $\bar{L}^{r}$ by $\bar{L}^{r}_{t}:=3\int_{0}^{t}L_{s}^{r}{\rm d}A_{s}$, for all $t\in\mathbb{R}_{+}$. Clearly, $\bar{L}^{r}$ is càdlàg, increasing and non-negative. Let $\lambda$ and $\beta$ be the functional and the increasing process associated to the compensated Poisson random measure $\tilde{\mu}$ as in [43, Theorem 31.9]. For an $\mathcal{H}$-valued càdlàg adapted process $Y$, we will use the abbreviations $a(t,Y)$, $b(t,Y)$ and $c(t,Y,x)$ for $a(t,\omega,Y_{t-}(\omega))$, $b(t,\omega,Y_{t-}(\omega))$ and $c(t,\omega,Y_{t-}(\omega),x)$, respectively. Moreover, we will use the shorthand notation $\bar{c}(t,Y)$ for the map $x\mapsto c(t,\omega,Y_{t-}(\omega),x)$. According to the Remark on [43, page 245], for all $\mathcal{H}$-valued càdlàg adapted processes $Y,Z$ and all $t\in\mathbb{R}_{+}$, it holds that
@@
By [43, Theorems 34.7 and 35.2], for each $\mathcal{F}_{0}$-measurable random variable $y_{0}:\Omega\to\mathcal{H}$ there exists a unique strong solution $Y$ to the SDE (4.4) with $Y_{0}=y_{0}$, where the stochastic integral with respect to the Brownian motion process $W$ is understood in the sense of [43, Chapter 26].
In order to complete the proof, it remains to prove that the stochastic integral with respect to $W$ coincides with the isometric stochastic integral. To this effect, note that the process $A$ given by $A_{t}:=2(t+1)$, for all $t\in\mathbb{R}_{+}$, is a control process of $W$ in the sense of [43, Definition 23.13]. Indeed, let $G$ be a separable Hilbert space, $\Phi$ an elementary $L(\mathbb{R}^{d},G)$-valued predictable process and $\tau$ a stopping time. Then, by Doob’s $L^{2}$-inequality and the Itô isometry, we have that
@@
In view of [43, Remark 26.2], this suffices to complete the proof. ∎
We shall now make use of Theorem 4.1 in order to prove a general existence and uniqueness result for SPDEs with random locally Lipschitz coefficients. Let $H$ be a separable Hilbert space and $A$ the generator of a $C_{0}$-semigroup $(S_{t})_{t\geq 0}$ on $H$. Furthermore, let
@@
$X_{t}=S_{t}x_{0}+\int_{0}^{t}S_{t-s}\alpha(s,X_{s}){\rm d}s+\int_{0}^{t}S_{t-s}\beta(s,X_{s}){\rm d}W_{s}+\int_{0}^{t}S_{t-s}\gamma(s,X_{s-},x)\tilde{\mu}({\rm d}s,{\rm d}x).$
@@
Suppose that the semigroup $(S_{t})_{t\geq 0}$ is pseudo-contractive, i.e., there exists a constant $\eta\geq 0$ such that
(4.12) $\|S_{t}\|\leq e^{\eta t},\quad\text{for all $t\geq 0$.}$
@@
Suppose that there exists a constant $\eta\geq 0$ such that (4.12) holds. Then there exist another separable Hilbert space $\mathcal{H}$, a $C_{0}$-group $(U_{t})_{t\in\mathbb{R}}$ on $\mathcal{H}$ satisfying
(4.19) $\|U_{t}\|=e^{\eta t},\quad\text{for all $t\in\mathbb{R}$,}$
@@
$\pi U_{t}\ell=S_{t},\quad\text{for all $t\in\mathbb{R}_{+}$,}$
@@
(4.20) $a(t,\omega,y)$ $:=U_{-t}\ell\alpha(t,\omega,\pi U_{t}y),$
(4.21) $b(t,\omega,y)$ $:=U_{-t}\ell\beta(t,\omega,\pi U_{t}y),$
(4.22) $c(t,\omega,y,x)$ $:=U_{-t}\ell\gamma(t,\omega,\pi U_{t}y,x).$
Note that $a$ and $b$ are $\mathcal{P}\otimes\mathcal{B}(\mathcal{H})$-measurable and $c$ is $\mathcal{P}\otimes\mathcal{B}(\mathcal{H})\otimes\mathcal{B}(E)$-measurable. This follows because the projection $(t,\omega)\mapsto t$ is $\mathcal{P}/\mathcal{B}(\mathbb{R}_{+})$-measurable and the mapping $(t,y)\mapsto U_{t}y$ is continuous.
@@
$\|a(t,\omega,y)-a(t,\omega,z)\|_{\mathcal{H}}^{2}$ $=e^{-2\eta t}\|\alpha(t,\omega,\pi U_{t}y)-\alpha(t,\omega,\pi U_{t}z)\|_{H}^{2}$
$\leq e^{-2\eta t}\bar{L}_{t}^{r}(\omega)\|\pi U_{t}y-\pi U_{t}z\|_{H}^{2}\leq\bar{L}_{t}^{r}(\omega)\|y-z\|_{\mathcal{H}}^{2},$
@@
$\|a(t,\omega,y)\|_{\mathcal{H}}^{2}$ $=e^{-2\eta t}\|\alpha(t,\omega,\pi U_{t}y)\|_{H}^{2}\leq e^{-2\eta t}L_{t}(\omega)(1+\|\pi U_{t}y\|_{H}^{2})$
@@ 4.2. Well-posedness of the real-world HJMM SPDE
(4.23) $\eta_{t}(\xi)=S_{t}\eta_{0}(\xi)+\int_{0}^{t}S_{t-s}\alpha(s,s+\xi){\rm d}s+\int_{0}^{t}S_{t-s}\beta(s,s+\xi){\rm d}W_{s}+\int_{0}^{t}\int_{E}S_{t-s}\gamma(s,s+\xi,x)\tilde{\mu}({\rm d}s,{\rm d}x),$
for all $t\geq 0$, where $(S_{t})_{t\geq 0}$ is the shift semigroup acting on the second time argument of $\alpha$, $\beta$ and $\gamma$. As in Section 4.1, we work under the standing assumption that $\mu({\rm d}t,{\rm d}x)$ is a homogeneous Poisson random measure with compensator $\nu({\rm d}t,{\rm d}x)=F({\rm d}x){\rm d}t$ and $E$ is a locally compact space.
@@
$V_{K}:\mathbb{R}_{+}\to\mathbb{R}_{+},\quad V_{K}(r):=r(1+r)\exp(Kr),$
and we introduce the strictly increasing function
$w_{K}:\mathbb{R}_{+}\to\mathbb{R}_{+},\quad w_{K}(r):=W_{K}(r)\wedge r.$
As usual, for $p\in(0,+\infty)$, we denote by $L^{p}(F)$ the space of all measurable functions $f:E\to\mathbb{R}$ such that $\int_{E}|f(x)|^{p}F({\rm d}x)<+\infty$.
@@
There exist $\rho^{\prime}>\rho$, an optional locally bounded non-negative process $\Lambda$, a non-negative function $\kappa\in L^{1}(F)\cap L^{2}(F)\cap L^{3}(F)$, a constant $M_{\beta}\in\mathbb{R}_{+}$ and an increasing function $M_{\gamma}:\mathbb{R}_{+}\to\mathbb{R}_{+}$ such that the following hold:
@@
    the processes $\lambda:\Omega\times\mathbb{R}_{+}\to\mathbb{R}^{d}$ and $b:\Omega\times\mathbb{R}_{+}\to\mathbb{R}^{(m+1)\times d}$ are predictable and locally bounded.
@@
    (4.28) $\big{|}\big{(}1+\psi_{t}(\omega,x)\big{)}\big{(}1+c_{t}^{i}(\omega,x)\big{)}\big{|}\leq\Lambda_{t}(\omega)\kappa(x),\qquad\text{ for all }(\omega,t,x)\in\Omega\times\mathbb{R}_{+}\times E.$
@@
    (4.30) $\|\beta(h)\|_{\rho,(m+1)\times d}$ $\leq M_{\beta}\sqrt{1+\|h\|_{\rho,m+1}},\qquad\text{ for all }h\in H^{m+1}_{\rho}.$
@@
    The function $\gamma:H_{\rho}^{m+1}\times E\to H_{\rho}^{m+1}$ is $\mathcal{B}(H_{\rho}^{m+1})\otimes\mathcal{B}(E)$-measurable and satisfies
    (4.31) $\gamma(\cdot,x)\in{\rm Lip}^{{\rm loc}}(H^{m+1}_{\rho},H^{0,m+1}_{\rho^{\prime}}),\qquad\text{for all $x\in E$.}$
    Furthermore, for each $r\in\mathbb{R}_{+}$ and $h,g\in H_{\rho}^{m+1}$ with $\|h\|_{\rho,m+1}\vee\|g\|_{\rho,m+1}\leq r$, it holds that
@@
    (4.33) $\|\gamma(h,x)\|_{\rho^{\prime},m+1}$ $\leq w_{K_{\rho,\rho^{\prime}}}\big{(}\kappa(x)(1+\|h\|_{\rho,m+1})\big{)},\qquad\text{ for all }(h,x)\in H_{\rho}^{m+1}\times E.$
@@
Beyond the extension to a multi-dimensional setup, Assumption 4.10 significantly weakens the requirements of [25] by replacing global Lipschitz continuity and the boundedness condition imposed in [25, Assumption 3.1] with local Lipschitz continuity and a growth condition, respectively. Besides its mathematical interest, this generalization is essential in our framework in order to consider models that might not admit a risk-neutral probability. Indeed, translating the technical requirements of [25] onto the real-world probability would lead to restrictive conditions on the market prices of risk which would typically imply the true martingale property of an LMD and, hence, the existence of a risk-neutral probability, thus precluding models for which the latter fails to exist.
@@
$\alpha^{i}_{1}(h)$ $:=\sum_{j=1}^{d}\beta^{i,j}(h)\bar{\beta}^{i,j}(h),$
$\alpha^{i}_{2}(t,\omega,h)$ $:=\sum_{j=1}^{d}\bigl{(}\lambda^{j}_{t}(\omega)+b^{i,j}_{t}(\omega)\bigr{)}\beta^{i,j}(h),$
$\alpha^{i}_{3}(t,\omega,h)$ $:=\int_{E}\gamma^{i}(h,x)\Bigl{(}1-e^{-\bar{\gamma}^{i}(h,x)}\bigl{(}1+\psi_{t}(\omega,x)\bigr{)}\bigl{(}1+c^{i}_{t}(\omega,x)\bigr{)}\Bigr{)}F({\rm d}x).$

### v3

@@ 3.5. An example: the minimal market model
We present a simple example of a model which falls into the real-world HJM setup developed above, while not admitting a risk-neutral probability. The example corresponds to an extension of the minimal market model (MMM), see [51, Chapter 13] and [13, Section 3.2]. Denoting by $X^{0}=\exp(\int_{0}^{\cdot}r_{s}{\rm d}s)$ the numéraire, we recall that in the stylized MMM the $X^{0}$-discounted growth-optimal portfolio (GOP) process $\bar{X}^{*}=(\bar{X}^{*}_{t})_{t\geq 0}$ satisfies the SDE
@@
with a one-dimensional Brownian process $W$ and where the drift $\alpha^{*}$ is modeled as a function of the form $\alpha^{*}(t)=\alpha_{0}\exp(\eta t)$, for all $t\geq 0$. The model is parameterized by a positive initial value $\alpha_{0}>0$ and a constant net growth rate $\eta>0$. We let $\{S^{i};i\in I\}$ be a family of strictly positive processes representing the spot processes. Similarly to [49, Section 3], we impose the following assumption. [^4] We recall that two processes $X=(X_{t})_{t\geq 0}$ and $Y=(Y_{t})_{t\geq 0}$ on $(\Omega,\mathbb{F},\mathbb{P})$ are said to be $\mathbb{F}$-conditionally independent if $\mathbb{E}[f(X_{T})g(Y_{T})|\mathcal{F}_{t}]=\mathbb{E}[f(X_{T})|\mathcal{F}_{t}]\mathbb{E}[g(Y_{T})|\mathcal{F}_{t}],\qquad\text{ for all }0\leq t\leq T<+\infty,$ for all measurable functions $f,g:\mathbb{R}\to\mathbb{R}_{+}$. This property is obviously satisfied if one of the two processes is deterministic.
@@
It is well-known that the MMM does not admit a risk-neutral probability (see [51, Section 13.3]). Therefore, prices are computed by relying on the real-world pricing formula (see [51, Section 10.4]). In particular, the price $B^{0}(t,T)$ of a riskless bond is given by
@@
The example described in this section illustrates a generic way of constructing real-world models satisfying market viability, without requiring risk-neutral probabilities. One can start from a strictly positive local martingale $D$, a numéraire $X^{0}$ and a family $\{S^{i}:i\in I\}$ of strictly positive spot processes. Generalizing (3.23) and (3.25), one can then specify riskless and risky bond prices as
(3.26) $B^{i}(t,T):=\mathbb{E}\left[\frac{D_{T}}{D_{t}}\frac{X^{0}_{t}}{X^{0}_{T}}\frac{S^{i}_{T}}{S^{i}_{t}}\bigg|\mathcal{F}_{t}\right],\qquad\text{ for all }0\leq t\leq T<+\infty\text{ and }i\in I_{0},$
where $S^{0}\equiv 1$. By construction, $D$ is a LMD and, therefore, the conditions of Theorem 3.4 are satisfied. Moreover, all term structures are fairly priced by $D$, in the sense of Definition 3.8.
@@ 4.1. A general existence and uniqueness result for SPDEs
Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered probability space as introduced in Section 3.1 and supporting a $d$-dimensional Brownian motion $W$ and a homogeneous Poisson random measure $\mu({\rm d}t,{\rm d}x)$ on $\mathbb{R}_{+}\times E$ with compensator $\nu({\rm d}t,{\rm d}x)=F({\rm d}x){\rm d}t$, where $E$ is a locally compact space equipped with its Borel sigma-field $\mathcal{B}(E)$ and $F$ is a sigma-finite measure on $(E,\mathcal{B}(E))$. The compensated random measure is denoted by $\tilde{\mu}({\rm d}t,{\rm d}x):=\mu({\rm d}t,{\rm d}x)-F({\rm d}x){\rm d}t$.
@@
Let $\overline{W}$ be the $\mathbb{R}^{d+1}$-valued semimartingale given by $\overline{W}_{t}:=(t,W_{t})$, for all $t\in\mathbb{R}_{+}$. Let $r\in\mathbb{R}_{+}$ be arbitrary and let $A$ be a control process for $\overline{W}$ in the sense of [48, Definition 23.13], which exists by [48, Theorem 23.11]. Since $L^{r}$ is locally bounded, we can define the process $\bar{L}^{r}$ by $\bar{L}^{r}_{t}:=3\int_{0}^{t}L_{s}^{r}{\rm d}A_{s}$, for all $t\in\mathbb{R}_{+}$. Clearly, $\bar{L}^{r}$ is càdlàg, increasing and non-negative. Let $\lambda$ and $\beta$ be the functional and the increasing process associated to the compensated Poisson random measure $\tilde{\mu}$ as in [48, Theorem 31.9]. For an $\mathcal{H}$-valued càdlàg adapted process $Y$, we will use the abbreviations $a(t,Y)$, $b(t,Y)$ and $c(t,Y,x)$ for $a(t,\omega,Y_{t-}(\omega))$, $b(t,\omega,Y_{t-}(\omega))$ and $c(t,\omega,Y_{t-}(\omega),x)$, respectively. Moreover, we will use the shorthand notation $\bar{c}(t,Y)$ for the map $x\mapsto c(t,\omega,Y_{t-}(\omega),x)$. According to the Remark on [48, page 245], for all $\mathcal{H}$-valued càdlàg adapted processes $Y,Z$ and all $t\in\mathbb{R}_{+}$, it holds that
@@
By [48, Theorems 34.7 and 35.2], for each $\mathcal{F}_{0}$-measurable random variable $y_{0}:\Omega\to\mathcal{H}$ there exists a unique strong solution $Y$ to the SDE (4.4) with $Y_{0}=y_{0}$, where the stochastic integral with respect to the Brownian motion process $W$ is understood in the sense of [48, Chapter 26].
In order to complete the proof, it remains to prove that the stochastic integral with respect to $W$ coincides with the isometric stochastic integral. To this effect, note that the process $A$ given by $A_{t}:=2(t+1)$, for all $t\in\mathbb{R}_{+}$, is a control process of $W$ in the sense of [48, Definition 23.13]. Indeed, let $G$ be a separable Hilbert space, $\Phi$ an elementary $L(\mathbb{R}^{d},G)$-valued predictable process and $\tau$ a stopping time. Then, by Doob’s $L^{2}$-inequality and the Itô isometry, we have that
@@
In view of [48, Remark 26.2], this suffices to complete the proof. ∎
We shall now make use of Theorem 4.1 in order to prove a general existence and uniqueness result for SPDEs with random locally Lipschitz coefficients. Let $H$ be a separable Hilbert space and $A$ the generator of a $C_{0}$-semigroup $(\mathcal{S}_{t})_{t\geq 0}$ on $H$. Furthermore, let
@@
$X_{t}=\mathcal{S}_{t}x_{0}+\int_{0}^{t}\mathcal{S}_{t-s}\alpha(s,X_{s}){\rm d}s+\int_{0}^{t}\mathcal{S}_{t-s}\beta(s,X_{s}){\rm d}W_{s}+\int_{0}^{t}\mathcal{S}_{t-s}\gamma(s,X_{s-},x)\tilde{\mu}({\rm d}s,{\rm d}x).$
@@
Suppose that the semigroup $(\mathcal{S}_{t})_{t\geq 0}$ is pseudo-contractive, i.e., there exists a constant $\eta\geq 0$ such that
(4.12) $\|\mathcal{S}_{t}\|\leq e^{\eta t},\quad\text{for all $t\geq 0$.}$
@@
Suppose that there exists a constant $\eta\geq 0$ such that (4.12) holds. Then there exist another separable Hilbert space $\mathcal{H}$, a $C_{0}$-group $(\mathcal{U}_{t})_{t\in\mathbb{R}}$ on $\mathcal{H}$ satisfying
(4.19) $\|\mathcal{U}_{t}\|=e^{\eta t},\quad\text{for all $t\in\mathbb{R}$,}$
@@
$\pi\,\mathcal{U}_{t}\ell=\mathcal{S}_{t},\quad\text{for all $t\in\mathbb{R}_{+}$,}$
@@
(4.20) $a(t,\omega,y)$ $:=\mathcal{U}_{-t}\ell\alpha(t,\omega,\pi\mathcal{U}_{t}y),$
(4.21) $b(t,\omega,y)$ $:=\mathcal{U}_{-t}\ell\beta(t,\omega,\pi\mathcal{U}_{t}y),$
(4.22) $c(t,\omega,y,x)$ $:=\mathcal{U}_{-t}\ell\gamma(t,\omega,\pi\mathcal{U}_{t}y,x).$
Note that $a$ and $b$ are $\mathcal{P}\otimes\mathcal{B}(\mathcal{H})$-measurable and $c$ is $\mathcal{P}\otimes\mathcal{B}(\mathcal{H})\otimes\mathcal{B}(E)$-measurable. This follows because the projection $(t,\omega)\mapsto t$ is $\mathcal{P}/\mathcal{B}(\mathbb{R}_{+})$-measurable and the mapping $(t,y)\mapsto\mathcal{U}_{t}y$ is continuous.
@@
$\|a(t,\omega,y)-a(t,\omega,z)\|_{\mathcal{H}}^{2}$ $=e^{-2\eta t}\|\alpha(t,\omega,\pi\mathcal{U}_{t}y)-\alpha(t,\omega,\pi\mathcal{U}_{t}z)\|_{H}^{2}$
$\leq e^{-2\eta t}\bar{L}_{t}^{r}(\omega)\|\pi\mathcal{U}_{t}y-\pi\mathcal{U}_{t}z\|_{H}^{2}\leq\bar{L}_{t}^{r}(\omega)\|y-z\|_{\mathcal{H}}^{2},$
@@
$\|a(t,\omega,y)\|_{\mathcal{H}}^{2}$ $=e^{-2\eta t}\|\alpha(t,\omega,\pi\mathcal{U}_{t}y)\|_{H}^{2}\leq e^{-2\eta t}L_{t}(\omega)(1+\|\pi\mathcal{U}_{t}y\|_{H}^{2})$
@@ 4.2. Well-posedness of the real-world HJMM SPDE
(4.23) $\eta_{t}(\xi)=\mathcal{S}_{t}\eta_{0}(\xi)+\int_{0}^{t}\mathcal{S}_{t-s}\alpha(s,s+\xi){\rm d}s+\int_{0}^{t}\mathcal{S}_{t-s}\beta(s,s+\xi){\rm d}W_{s}+\int_{0}^{t}\int_{E}\mathcal{S}_{t-s}\gamma(s,s+\xi,x)\tilde{\mu}({\rm d}s,{\rm d}x),$
for all $t\geq 0$, where $(\mathcal{S}_{t})_{t\geq 0}$ is the shift semigroup acting on the second time argument of $\alpha$, $\beta$ and $\gamma$, so that $\mathcal{S}_{t}\alpha(s,\cdot)=\alpha(s,t+\cdot)$, and similarly for the other terms appearing in (4.23). As in Section 4.1, we work under the standing assumption that $\mu({\rm d}t,{\rm d}x)$ is a homogeneous Poisson random measure with compensator $\nu({\rm d}t,{\rm d}x)=F({\rm d}x){\rm d}t$ and $E$ is a locally compact space.
@@
(4.28) $V_{K}:\mathbb{R}_{+}\to\mathbb{R}_{+},\quad V_{K}(r):=r(1+r)\exp(Kr).$
It can be easily seen that $W_{K}(r)\leq r$, for all $r\in\mathbb{R}_{+}$. As usual, for $p\in(0,+\infty)$, we denote by $L^{p}(F)$ the space of all measurable functions $f:E\to\mathbb{R}$ such that $\int_{E}|f(x)|^{p}F({\rm d}x)<+\infty$.
@@
There exist $\rho^{\prime}>\rho$, an optional locally bounded non-negative process $\Lambda$, a non-negative function $\kappa\in L^{1}(F)\cap L^{2}(F)\cap L^{3}(F)$, a constant $M_{\beta}\in\mathbb{R}_{+}$, an increasing function $M_{\gamma}:\mathbb{R}_{+}\to\mathbb{R}_{+}$ and a constant $N_{\gamma}\in\mathbb{R}_{+}$ such that the following hold:
@@
    the processes $\lambda:\Omega\times\mathbb{R}_{+}\to\mathbb{R}^{d}$ and $b:\Omega\times\mathbb{R}_{+}\to\mathbb{R}^{(m+1)\times d}$ are predictable and locally bounded;
@@
    (4.29) $\big|\big(1+\psi_{t}(\omega,x)\big)\big(1+c_{t}^{i}(\omega,x)\big)\big|\leq\Lambda_{t}(\omega)\kappa(x),\qquad\text{ for all }(\omega,t,x)\in\Omega\times\mathbb{R}_{+}\times E;$
@@
    (4.31) $\|\beta(h)\|_{\rho,(m+1)\times d}$ $\leq M_{\beta}\sqrt{1+\|h\|_{\rho,m+1}},\qquad\text{ for all }h\in H^{m+1}_{\rho};$
@@
    the function $\gamma:H_{\rho}^{m+1}\times E\to H_{\rho}^{m+1}$ is $\mathcal{B}(H_{\rho}^{m+1})\otimes\mathcal{B}(E)$-measurable and, for each $r\in\mathbb{R}_{+}$ and $h,g\in H_{\rho}^{m+1}$ with $\|h\|_{\rho,m+1}\vee\|g\|_{\rho,m+1}\leq r$, it holds that
@@
    (4.33) $\|\gamma(h,x)\|_{\rho^{\prime},m+1}$ $\leq W_{K_{\rho,\rho^{\prime}}}\big(N_{\gamma}\kappa(x)(1+\|h\|_{\rho,m+1})\big),\qquad\text{ for all }(h,x)\in H_{\rho}^{m+1}\times E.$
@@
Note that condition (4.32) implies that
$\gamma(\cdot,x)\in{\rm Lip}^{{\rm loc}}(H^{m+1}_{\rho},H^{0,m+1}_{\rho^{\prime}}),\qquad\text{for all $x\in E$.}$
Remark 4.12.
Beyond the extension to a multi-dimensional setup, Assumption 4.10 significantly weakens the requirements of [25] by replacing global Lipschitz continuity and the boundedness condition imposed in [25, Assumption 3.1] with local Lipschitz continuity and a growth condition, respectively. In particular, we explicitly allow for unbounded coefficients $\beta$ and $\gamma$. Example 4.15 provides a model with unbounded $\beta$ and $\gamma$ for which Assumption 4.10 is fulfilled (and, therefore, Theorem 4.14 applies). This example covers the classical risk-neutral setup and thus demonstrates that our results strictly extend those of [25], even within the traditional risk-neutral setup. Besides its mathematical interest, this generalization is essential to accommodate models that may fail to admit a risk-neutral probability. Indeed, imposing the requirements of [25] under the real-world measure would typically translate into restrictive assumptions on the market prices of risk, often forcing the true martingale property of an LMD and, hence, the existence of a risk-neutral measure. Our framework avoids this structural limitation while still covering the classical risk-neutral setup as a special instance.
@@
$\alpha^{i}_{1}(h)$ $=\sum_{j=1}^{d}\beta^{i,j}(h)\bar{\beta}^{i,j}(h),$
$\alpha^{i}_{2}(t,\omega,h)$ $=\sum_{j=1}^{d}\bigl(\lambda^{j}_{t}(\omega)+b^{i,j}_{t}(\omega)\bigr)\beta^{i,j}(h),$
$\alpha^{i}_{3}(t,\omega,h)$ $=\int_{E}\gamma^{i}(h,x)\Bigl(1-e^{-\bar{\gamma}^{i}(h,x)}\bigl(1+\psi_{t}(\omega,x)\bigr)\bigl(1+c^{i}_{t}(\omega,x)\bigr)\Bigr)F({\rm d}x).$